OpenStudy (anonymous):

Could someone walk me through this: 2^(x+1) = 15 Here's what I've got: 1) (x+1)log2 = log15 2) x+1 = log(15)/log(2) Am i doing it right?

OpenStudy (solomonzelman):

\[2^{(x+1)}=15\] Hit both sides with the log.

OpenStudy (solomonzelman):

I got disconnected, sorry.

OpenStudy (cggurumanjunath):

u r right ! @pabloecortez07

OpenStudy (cggurumanjunath):

x=(log(15)/log(2))-1

OpenStudy (solomonzelman):

\[(x+1)Log2=Log15\] Divide both sides by Log2 \[x+1=\frac{Log15}{Log2}\] \[x=\frac{Log15}{Log2}-1\]

OpenStudy (solomonzelman):

yes, as @CGGURUMANJUNATH, pointed out, you didn't finish, you didn't subtract 1 from both sides.

OpenStudy (solomonzelman):

this would help you more in the future, http://www.rapidtables.com/math/algebra/Ln.htm#ln rules

OpenStudy (anonymous):

Thanks to both of you.

OpenStudy (solomonzelman):

\(\huge {\color{blue} {anytime!} }\)